OpenStudy (anonymous):

Which expression is the first step you take when finding the quotient of five over six division symbol two over three? five-sixths times three-halves sififths times two-thirds two-thirds times five-sixths sififths times thee-halves

OpenStudy (johnweldon1993):

The first step to take...would be flip the fraction on the bottom and change the operation to multiplication...that is...?

OpenStudy (anonymous):

@johnweldon1993 it needs to be these 4 on the list

OpenStudy (anonymous):

five-sixths times three-halves sififths times two-thirds two-thirds times five-sixths sififths times thee-halves

OpenStudy (johnweldon1993):

And it IS one of those choices...read what was written and tell me what choice that represents

OpenStudy (anonymous):

is it a @johnweldon1993

OpenStudy (johnweldon1993):

That is correct!

OpenStudy (anonymous):

can i ask a few more

OpenStudy (anonymous):

if yes here is a problem What is the sum of three and three over eight plus two and three over eight in simplest form? six and three over eight six and three over four five and three over eight five and three over four

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

As in mixed numbers?

OpenStudy (anonymous):

yes i believe s0

OpenStudy (johnweldon1993):

\[3\frac{3}{8} + 2\frac{3}{8}\] Do you know how to convert these to improper fractions?

OpenStudy (anonymous):

i think so refresh me

OpenStudy (johnweldon1993):

1) Multiply the denominator and the whole number 2) Add the numerator to the result 3) place the result over the original denominator

OpenStudy (anonymous):

ok

OpenStudy (johnweldon1993):

So \[\large 3\frac{3}{8} = ?\] and \[\large2\frac{3}{8} = ?\]

OpenStudy (anonymous):

nevermind this is to complicated

OpenStudy (johnweldon1993):

\[\large 3\frac{3}{8} =\frac{(8 \times 3)+3}{8} = \frac{27}{8}\] \[\large 2\frac{3}{8} =\frac{(8 \times 2)+3}{8} = \frac{19}{8}\] \[\large \frac{27 + 19}{8} = \frac{46}{8} = \frac{23}{4}\] Turning this back into a mixed number is easy too...just divide....how many times does 4 go into 23? 5 times....that is 5 * 4 = 20 ...so leaves a remainder of 3..and you put that 3 over the original denominator again 5 and three over four

OpenStudy (anonymous):

What is the quotient of two and five over six divided by two over three in simplest form? one and eight over nine four and one over four three and one over four two and five over nine

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@happinessbreaksbones

OpenStudy (anonymous):

do u have anything yet